OpenStudy (anonymous):

Simplify tan(x)^2-tan(x)^2sin(x)^2

OpenStudy (anonymous):

Hi, the first step would be to factor out what the terms have in common.

OpenStudy (anonymous):

Do you see what they have in common?

OpenStudy (anonymous):

tan(x)^2 and - tan(x)^2sin(x)^2 both have tan(x)^2 in common

OpenStudy (anonymous):

tan(x)^2(1 - sin(x)^2)

OpenStudy (jdoe0001):

\(\bf sin^2(\theta)+cos^2(\theta)=1\implies cos^2(\theta)={\color{blue}{ 1-sin^2(\theta)}}\qquad tan(\theta)=\cfrac{sin(\theta)}{cos(\theta)}\\ \quad \\ \quad \\ tan(x)^2-tan(x)^2sin(x)^2\implies tan^2(x)[{\color{blue}{ 1-sin^2(x}})]\\ \quad \\ \cfrac{sin(\theta)}{cos(\theta)}[{\color{blue}{ 1-sin^2(x}})]\)

OpenStudy (anonymous):

The next step is to solve the identity sin(x)^2 + cos(x)^2 = 1 for 1 - sin(x)^2

OpenStudy (anonymous):

We can do this by subtracting sin(x)^2 from both sides.

OpenStudy (jdoe0001):

hmmm actually

OpenStudy (anonymous):

sin(x)^2 + cos(x)^2 - sin(x)^2 = 1 - sin(x)^2

OpenStudy (jdoe0001):

\(\bf sin^2(\theta)+cos^2(\theta)=1\implies cos^2(\theta)={\color{blue}{ 1-sin^2(\theta)}}\qquad tan(\theta)=\cfrac{sin(\theta)}{cos(\theta)}\\ \quad \\ \quad \\ tan(x)^2-tan(x)^2sin(x)^2\implies tan^2(x)[{\color{blue}{ 1-sin^2(x}})]\\ \quad \\ \cfrac{sin^2(\theta)}{cos^2(\theta)}[{\color{blue}{ 1-sin^2(x}})]\) better

OpenStudy (anonymous):

cos(x)^2 = 1 - sin(x)^2

OpenStudy (anonymous):

We can then substitute the 1 - sin(x)^2 with cos(x)^2

OpenStudy (anonymous):

tan(x)^2(1 - sin(x)^2) = tan(x)^2(cos(x)^2)

OpenStudy (anonymous):

Since tan(x) = sin(x)/cos(x) and tan(x)^2 = sin(x)^2/cos(x)^2, we can substitute again.

OpenStudy (anonymous):

tan(x)^2(cos(x)^2) = (sin(x)^2/cos(x)^2)(cos(x)^2)

OpenStudy (anonymous):

The cos(x)^2 cancels in the numerator and the denominator.

OpenStudy (anonymous):

We are left with sin(x)^2

OpenStudy (anonymous):

Thank you I understand what you are saying